OpenStudy (anonymous):

What is the standard form of the equation of the conic given by 2x^2 + 2y^2 + 4x - 12y - 22 = 0 ?

OpenStudy (anonymous):

https://people.richland.edu/james/lecture/m116/conics/conics.html consult the page..:D

OpenStudy (anonymous):

Still don't understand

OpenStudy (anonymous):

Ax2 + Bxy + Cy2 + Dx + Ey + F = 0 so just take a 2 common from your equation and then divide both sides of the equation with 2 :D

OpenStudy (anonymous):

Can you please walk me through the first steps?

OpenStudy (anonymous):

lemme do it for you and sorry for leaving..! lets take out 2 common 2x^2 + 2y^2 + 4x - 12y - 22 = 0 => 2 ( x^2 + y^2 + 2x - 6y - 11) = 0 now lets divide both sides with 2 => x^2 + y^2 + 2x - 6y - 11 = 0